Belle is 1/7 as old as Anne. In eight years, Belle will be 1/3 as old as Anne. Calculate their ages.
Thank you.

Answer by nerdybill(7384)

Belle is 1/7 as old as Anne. In eight years, Belle will be 1/3 as old as Anne. Calculate their ages.
.
Let A = Anne's age
then
(1/7)A = Belle's age
.
(1/7)A + 8 = (1/3)(A+8)
Multiplying both sides by 21 (to rid ourselves of the denominator) we get:
3A + 168 = 7(A+8)
3A + 168 = 7A + 56
168 = 4A + 56
112 = 4A
28 = A (Anne's age)
(1/7)A = 4 (Belle's age)